Steep safety ramps are built beside mountain highways to enable vehicles with defective brakes to stop safely. A truck enters a 250-m ramp at a high speed v0, and travels 180 m in 6 s at constant deceleration before its speed is reduced to v0/2. Assuming the same constant deceleration, determine (a) the additional time required for the truck to stop, (b) the additional distance traveled by the truck.



1
Expert's answer
2022-04-26T14:55:52-0400

"\frac v2=v-at,"

"s=vt-\frac{at^2}2," "\implies"

"a=\frac{3v^2}{8s},"


"0=\frac v2-at_1,"

"t_1=\frac{4s}{3v},"

"s_1=\frac v2 t_1-\frac{at_1^2}2=\frac s3=60~m,"

"t=\frac v{2a}=\frac{4s}{3v},"

"t_1=t=6~s."